OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the trig function; tan(105*) Please show detailed work, thank you...

OpenStudy (anonymous):

tan105=(tan135-tan30)/(1+tan135tan30)

OpenStudy (anonymous):

I've gotten to the point above, but I do not know how to simplify.

OpenStudy (anonymous):

(-1-sqrt3/3)/(1+(-1)(sqrt3/3))

OpenStudy (amistre64):

105 comes from what angles?

OpenStudy (anonymous):

135-30

OpenStudy (amistre64):

and 135 is also known as?

OpenStudy (amistre64):

180-135 = 45

OpenStudy (anonymous):

(3pi)/4

OpenStudy (amistre64):

45 is simpler to evaluate with

OpenStudy (amistre64):

i see you went straight for the tan(a+b) instead of wasting time with the sin(a+b)/cos(a+b)

OpenStudy (amistre64):

i can never really recall which tan is the right one so I just go by the sin cos sin(a+b) = sin(a)cos(b) + sin(b)cos(a) ------------------------------- cos(a+b) = cos(a)cos(b) -sin(a)sin(b) sin(30)cos(45) + sin(45)cos(30) ---------------------------- cos(30)cos(45) -sin(30)sin(45)

OpenStudy (amistre64):

1/2 sqrt(2)/2 + sqrt(2)/2 sqrt(3)/2 ------------------------------ sqrt(3)/2 sqrt(2)/2 - 1/2 sqrt(2)/2 sqrt(2)/4 + sqrt(6)/4 ------------------ sqrt(6)/4 - sqrt(2)/4 sqrt(2) + sqrt(6) ------------- sqrt(6) - sqrt(2) seems to fit

OpenStudy (amistre64):

do a conjugate of the bottom to see if we can simplify it more

OpenStudy (anonymous):

(-sqrt2+sqrt6)-(sqrt6-sqrt2)

OpenStudy (anonymous):

I do not know what a conjugate is

OpenStudy (amistre64):

sqrt(2) + sqrt(6) sqrt(6) + sqrt(2) ------------- -------------- sqrt(6) - sqrt(2) sqrt(6) + sqrt(2) 8 + 4 sqrt(3) ----------- 6 - 2 8 + 4 sqrt(3) ----------- 4 2 + sqrt(3) with any luck that checks out with the wolf

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=tan%28105%29 yep, just gotta place it in the correct quadrant for sign

OpenStudy (amistre64):

a conjugate is a name for a pair of ..... well (a+b) (a-b) is a conjugate pair

OpenStudy (anonymous):

OK, thank you for your time and effort

OpenStudy (amistre64):

youre welcome